All the major stock market indexes posted heavy losses in 2008. The mean one-year return for stocks in the S&P 500, a group of 500 very large companies, was -38.5%. The mean one year return for the NASDAQ, a group of 3,200 small and medium sized companies, was -40.5%. Historically, the one year return are approximately normal, the standard deviation in the S&P 500 is approximately 20%, and the standard deviation in the NASDAQ is approximately 30%. a. What is the probability that a stock in the S&P 500 gained value in 2008? b. What is the probability that a stock in the S&P 500 gained 10% more? c. What is the probability that a stock in the S&P 500 lost 50% more in 2008? d. What is the probability that a stock in the S&P 500 lost 60% more? e. Repeat (a) through (d) for a stock in the NASDAQ. f. Write a short summary on your findings. Be sure to include a discussion of the risks associated with a large standard deviation.
Answer by stanbon(75887)

All the major stock market indexes posted heavy losses in 2008. The mean one-year return for stocks in the S&P 500, a group of 500 very large companies, was -38.5%. The std = 20%
-----
The mean one year return for the NASDAQ, a group of 3,200 small and medium sized companies, was -40.5%. The std is 30%
------
a. What is the probability that a stock in the S&P 500 gained value in 2008?
z(0) = (0-(-38.5))/20 = 1.925
P(x > 0) = P(z > 1.925 = 0.0271
============================================
b. What is the probability that a stock in the S&P 500 gained 10% more?
Find the z-value of 10
Find the probability of z being >= that z-value
-------------------
c. What is the probability that a stock in the S&P 500 lost 50% more in 2008?
Find the z-value of -50
Find the probability of z being <= that z-value
--------------------------------
d. What is the probability that a stock in the S&P 500 lost 60% more?
? more than what?
e. Repeat (a) through (d) for a stock in the NASDAQ. f. Write a short summary on your findings. Be sure to include a discussion of the risks associated with a large standard deviation.
==============
Cheers,
Stan H.